OpenStudy (anonymous):

Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (aamirgsen):

x = -8/7, y = -4/7

OpenStudy (aamirgsen):

btw this should be in Mathematics :P

OpenStudy (anonymous):

what he said